OpenStudy (anonymous):

how do you write y = 2x2 + 6x + 4 in general form?

OpenStudy (anonymous):

do u know the general form?

OpenStudy (anonymous):

yes, y=a(x-h)^2+k

OpenStudy (anonymous):

did u need it in linear form?

OpenStudy (anonymous):

like the general form of a linear equation?

OpenStudy (anonymous):

just the general form of a quadratic equation

OpenStudy (anonymous):

okay, so u need to know the quadratic formula i believe..do u know it?

OpenStudy (anonymous):

general form> y=a(x-h)^2+k

OpenStudy (anonymous):

oh, yeah!

OpenStudy (anonymous):

quadratic formula> -b+-sqrt(b-4(a)(c))/2(a)

OpenStudy (anonymous):

but how does that help convert it to general form?

OpenStudy (anonymous):

I thought it just helped find the x-intercepts..

OpenStudy (anonymous):

if i can remember this correctly, i believe u need to first solve to get the values of a,b,c and then you just plug it into the general form... I'm not 100% sure tho... @abb0t, am i remembering this correctly? ;/

OpenStudy (anonymous):

I know how to convert it... the problem is the 2 in 2x^2

OpenStudy (abb0t):

General form of what?

OpenStudy (anonymous):

quadratic equations

OpenStudy (anonymous):

2 in 2x^2 is the problem.... you isolate the x variables getting: y-4=2x^2+6x then complete the square and balance the equation: 6/2=3 3^2=9, y-4+9=2x^2+6x+9, y+5=2x^2+6x+9 then convert the trinomial into a binomial... thus lies my dilemma

OpenStudy (anonymous):

sorry, i'm not very sure, cuz i don't think i remember this correctly ahha and i don't wanna help u wrongly haha @Mertsj probs knows :) good luck!!! :D

OpenStudy (anonymous):

thanks (:

OpenStudy (mertsj):

\[y=2x^2+6x+4\] \[y=2(x^2+3x+______ )+4\]

OpenStudy (mertsj):

Now complete the square by adding (3/2)^2 inside the parentheses: \[y=2(x^2+3x+(\frac{3}{2})^2)+4-\frac{9}{2}\]

OpenStudy (mertsj):

Now factor: \[y=2(x+\frac{3}{2})^2-\frac{1}{2}\]

OpenStudy (anonymous):

why 3/2?

OpenStudy (mertsj):

I want to show you something: \[x^2+6x+9=(x+3)^2\]

OpenStudy (mertsj):

\[x^2+8x+16=(x+4)^2\]

OpenStudy (anonymous):

I don't get what you did with: (3/2^2)+4−92

OpenStudy (mertsj):

Notice the relationship between the coefficient of x and the constant term in a trinomial square.

OpenStudy (mertsj):

If you take 1/2 the coefficient and square it, you get the constant term.

OpenStudy (mertsj):

So I took 1/2 of 3 and got 3/2. Then I squared it and added it to the x^2 + 3x to get a trinomial square.

OpenStudy (mertsj):

Now the (3/2)^2 was inside a parenthesis which has a 2 in front of it so I was really adding 2 times (3/2)^2 which is 2 times 9/4 which is 9/2. So since I could not change the equation, I then had to subtract 9/2.

OpenStudy (mertsj):

Follow me?

OpenStudy (anonymous):

I think so

OpenStudy (anonymous):

Okay, yes! I get what you did (: thank you! what were you trying to show me with those other two equations?

OpenStudy (mertsj):

The relationship between the coefficient of x and the constant term.

OpenStudy (anonymous):

oh, that half of the coefficient of x to the second power = the constant. okay, cool

OpenStudy (anonymous):

Thank you for helping me (:

OpenStudy (mertsj):

yw